# Passage-traversing optimal path planning with sampling-based algorithms

### v1

@@ 7.3.1 Planning efficiency with different obstacle numbers
Figure 22: Cost of the best path versus the number of valid samples in different planners and configuration spaces. The values are normalized so that the cost of the final optimal path is one.
@@
7.3.2 Planning efficiency with different sample numbers and cost evolution
Lastly, the planning time with different valid sample numbers is shown in Figure 21. The environment is populated with 40 obstacles. The planning time is averaged over five runs. Direct passage traversal PTOPP takes significantly more time than other planners as the sample number increases, while the proposed PTOPP dramatically reduces the time expense. In PRM-based methods, its performance is better or close to the benchmarks of clearance constraint and maximum clearance planning. In RRT, it consumes slightly more time than benchmarks only when the valid sample number gets extremely large. For planning time trends, RRT-based planners exhibit stronger linearity in the obstacle number, whereas the linearity of running time in the valid sample number is more evident in PRM-based planners. PRM-based planners also take shorter planning time when the sample number is large. Figure 22 illustrates the evolution of costs of the best paths during planning to highlight the fast convergence of PTOPP. Because MPW-PTOPP and maximum clearance planning find homotopic optimal paths in 2D space, they are compared together with the shortest path planning baseline. MPW-PTOPP converges to the optimal cost much faster than others due to the discreteness of its cost. Finite cost values exist in PTOPP and thus can be explored rapidly, as illustrated by stages in Figure 22 and 22. Clearance also gets updated periodically in RRT, but is less efficiently explored for its value continuity. PRM-based planners demonstrate faster convergence than RRT counterparts because random roadmaps can explore a much larger configuration space partition than random trees in RRT given the same number of samples. Random trees only grow from leaves in the sampling process. Note that the maximum clearance in PRM can decrease with increasing samples. This is attributed to that the near neighbor search radius $r_{PRM^{*}}$ is set to decrease with samples. As a consequence, samples of large clearance may not be reachable from path nodes given a small $r_{PRM^{*}}$. To address this, a large $r_{PRM^{*}}$ is required for dense samples at the cost of higher time consumption.
@@ 8 Discussion and conclusion
Passages provide advantages over clearance in accessible free space measurement. Many approaches to enlarging path clearance, such as path search in Voronoi graphs and path postprocessing, are unable to globally maximize clearance. Though maximum clearance planning can be employed, the degradation problem exists as in MPW-PTOPP. Since path clearance values are continuous, they cannot be optimized the same as discrete passage widths in PTOPP. Passage graphs offer a new perspective to organize objects potential in scene understanding. Narrow passage regions may be used to guide sampling within for efficient path finding. The Gabriel condition is proposed for passage detection because of its intuitive physical interpretation in robotics contexts. Another possible choice of proximity graph is the relative neighborhood graph (RNG), a subgraph of the Gabriel graph (Toussaint, 1980; Melchert, 2013). RNG only maintains the shortest edges of Delaunay triangles, imposing a stricter passage criterion than the Gabriel condition. The present pipeline of passage detection and cell decomposition still applies. Beyond the 2D and 3D workspace, the concept of passages is promising in higher dimensional spaces, e.g., the joint space, in describing confined configurations. But efficient characterization of subjects, e.g., obstacles, in high dimensional space is complicated (Dai et al., 2024).
Accessible free space optimization for paths is studied in this article as one most direct and important application of PTOPP. Other PTOPP problem instances exist since PTOPP has good configurability and applicability to formulate and solve various problems related to passages. For example, passages may be abstracted as toll stations that incur costs to pass. For good path smoothness, one may minimize the number of traversed passages. The principles of compatible path costs provide important guidance for cost design in OPP. The presented algorithmic pipeline of PTOPP is general and aligns with sampling-based optimal planners. Thus, common techniques to accelerate planning and improve the solution quality, such as bidirectional planning and biasing samples to the goal region, are applicable. Apart from sampling-based methods, pure cell-based path planning methods are enabled by Gabriel cells. They can reach a much faster planning speed with no need for a large number of random samples. Their dependent procedures, however, are covered in the proposed algorithms.
@@ A.1 Geodesic distance traversed in Delaunay graphs in passage detection
When extracting the Gabriel graph from a Delaunay graph, it suffices to check all the Delaunay graph edges if the graph nodes are simple points. For edge $(\mathbf{x}_{i},\mathbf{x}_{j})$, the graph nodes that possibly violate the Gabriel condition are also restricted to nodes adjacent to $\mathbf{x}_{i}$ or $\mathbf{x}_{j}$. When determining passage candidates from the Delaunay graph $\mathcal{DG}(C_{c})$ constructed from the obstacle centroids set $C_{c}$, solely examining edges of $\mathcal{DG}(C_{c})$ may miss some passages because obstacles are volumetric. The solution is to increase the check range defined by the geodesic distance, which is one in the point node case. The smallest geodesic distance $k_{gd}$ in $\mathcal{DG}(C_{c})$ is described in (9) and adopted as two in our implementation, which proves to be sufficient to find all passages under test conditions. The major factors influencing $k_{gd}$ include dimension and distribution properties of obstacles. The more obstacles can be treated as simple points (e.g., with small sizes and in large environments), the smaller $k_{gd}$ is.
Here we report data on detected passage numbers in various setups with $k_{gd}=1$ and $2$. It empirically verifies that a small $k_{gd}>1$ is sufficient to detect all passages in most cases. As demonstrated in Table 5 and 6, when obstacle sizes have limited variations, $k_{gd}=2$ reports all passages in all trials. This corresponds to the side length range of $[20,60]$ used in experiments. As expected, only checking edges of $\mathcal{DG}(C_{c})$, i.e., $k_{gd}=1$, fails to find all passages in most tests. If obstacle sizes have extreme variations, $k_{gd}=2$ may also fail. When the obstacle side length range is set as $[1,60]$, obstacle sizes may differ by thousands of times. Failure to detect all passages occurs occasionally for $k_{gd}=2$, especially when obstacles are dense. Importantly, only a negligible number of passages are missed in failure cases as listed in Table 6, which is acceptable in most practical scenes. To find all passages, $k_{gd}$ only needs to be larger than two. $k_{gd}$ does not change the detection complexity because the expected degree of a Delaunay graph node is $O(1)$ and at most six (Mark et al., 2008). The computational efficiency is guaranteed to outperform the brute-force check strategy.
Table 5: Number of trails that report all passages in different settings. For a given obstacle number and side length range, 30 trials of randomly generated obstacle distributions are conducted. The environment dimension is fixed as $1000\times 600$.
@@
A.2 Compatible path cost design in GPW-PTOPP
In general, the cost form in (20) is not order-preserving as a sum of paired weight and width products. The design objective is to make the order of path costs consistent with the reversed lexicographic order of passage width vectors by setting appropriate weights. The lexicographic order of passage vectors is preserved in path concatenation since the following relation holds
where $\sqcup$ stands for inserting a new entry into one sorted vector while maintaining the sorted order. This is easily verified by comparing the new element’s insertion positions in two vectors. $\mathbf{p}_{\sigma}$ is fixed sized in PTOPP and the insertion is accomplished by first inserting the new element and then truncating the last and also the largest element. To ensure the transition in (22) for path cost (20), weight values in $\mathbf{w}_{p}$ are analyzed recursively. Start with the basic case of 2-GPW-PTOPP, i.e., $\textit{k}=2$. Let $\mathbf{p}_{\sigma_{1}}=[p_{1,1},p_{1,2}]^{\mathsf{T}},\mathbf{p}_{\sigma_{2}}=[p_{2,1},p_{2,2}]^{\mathsf{T}},\mathbf{w}_{p}=[w_{1},w_{2}]^{\mathsf{T}}$ and $\mathbf{p}_{\sigma_{1}}\prec_{\text{lex}}\mathbf{p}_{\sigma_{2}}$. $f_{c}(\sigma_{1})>f_{c}(\sigma_{2})$ implies that
$w_{1}p_{1,1}+w_{2}p_{1,2}<w_{1}p_{2,1}+w_{2}p_{2,2}$ (32)
The case where $p_{1,1}=p_{2,1},p_{1,2}<p_{2,2}$ is trivial. If $p_{1,1}<p_{2,1}$, the full constraint on weights is expressed as
@@ A.2 Compatible path cost design in GPW-PTOPP
Extending to $\textit{k}=3$, the similar ratio constraint on weights is derived. If $p_{1,1}=p_{2,1}$, the constraint in (33) applies to $w_{2},w_{3}$ except that choices of passage widths decrease by one because $p_{1,1}=p_{2,1}$ is removed. But the scaled value in (34) is still applicable to set $w_{2}/w_{3}>\max\Delta p/\min\Delta p$. If $p_{1,1}<p_{2,1}$, the following inequality holds
@@
By this means, the reversed lexical orders of passage width vectors are maintained in (20). The ratios between contiguous weight items are required to be sufficiently large. In practice, any large ratio that meets (37) works. Note that if $\mathbf{p}_{\sigma_{1}},\mathbf{p}_{\sigma_{2}}$ take arbitrary values out from $\mathbb{P}$, (22) may be violated. In such cases, it implies there exists $\mathbf{p}_{\sigma_{1},k}\gg\mathbf{p}_{\sigma_{2},k}$ for some $k>j$ in (21), also representing a better $\mathbf{p}_{\sigma}$ candidate for admitting significantly wider passages. Hence, it suffices to use $w_{p,i}/w_{p,i+1}\gg 1$ in GPW-PTOPP.
There are other possible constructions of compatible GPW-PTOPP path costs, such as $f_{c}(\sigma)=-\sum p_{i}$, $p_{i}\in\mathbf{p}_{\sigma}$ and its combination with the path length $f_{c}=\text{Len}(\sigma)-\sum p_{i}$ as in (24). They have simple structures and good interpretability. But the problem is that while these costs favor wide passages, they are unable to prioritize avoiding narrow passages. For example, $\mathbf{p}_{\sigma}=[1,9]^{\mathsf{T}}$ leads to the same cost as $\mathbf{p}_{\sigma}=[5,5]^{\mathsf{T}}$ though it contains a much smaller passage width. For this reason, they are not adopted in this paper for accessible free space optimization purposes.

### v2

@@ 7.3.1 Planning efficiency with different obstacle numbers
7.3.2 Planning efficiency with different valid sample numbers and cost evolution
Planning time with varying valid sample numbers $|V|$ is shown in Figure 19. The obstacle number $m=40$ and the planning time is averaged over five runs. PTOPP using passage traversal takes significantly more time than other planners as $|V|$ increases, while the proposed PTOPP remarkably reduces the time expense. In PRM-based planners, the proposed method has planning time shorter than or similar to clearance constraint and maximum clearance planning. In RRT-based planners, it consumes slightly more time than benchmarks only when $|V|$ gets extremely large. As for planning time trends, RRT-based planners exhibit stronger linearity w.r.t. $m$, whereas PRM-based planners exhibit stronger linearity w.r.t. $|V|$. PRM-based planners run faster when $|V|$ gets large.
Figure 18: Average planning time with varying obstacle number $m$. Dashed and solid lines correspond to PRM-based and RRT-based algorithms, respectively. For each $m$, 30 trials in random obstacle distributions are conducted.
Figure 19: Average path planning time with varying valid sample number $|V|$. For each $|V|$, five trials in the same environment with 40 obstacles are conducted.
Lastly, Figure 20 illustrates the evolution of costs of the best paths during planning to highlight the fast convergence of PTOPP. Because MPW-PTOPP and maximum clearance planning find homotopic optimal paths in 2D space, they are compared along with the shortest path planning baseline. MPW-PTOPP converges to the optimal cost much faster than others due to the discreteness of its cost. Finite cost values exist in PTOPP and thus are explored quickly, as illustrated by stages in Figure 20 and 20. Clearance also gets updated periodically in RRT, but is less efficiently explored for its value continuity. PRM-based planners demonstrate faster convergence than RRT counterparts because random roadmaps can explore a much larger configuration space partition than random trees do in RRT given the same number of samples. Random trees only grow from leaves. Note that the maximum clearance in PRM may decrease with the sample number increasing. This counterintuitive phenomenon is attributed to that the near neighbor search radius $r_{PRM^{*}}$ is set to decrease with samples. Samples of large clearance may not be reachable from path nodes if $r_{PRM^{*}}$ is small.
Figure 20: Cost of the best path versus the number of valid samples $|V|$ in different planners and configuration spaces. The values are normalized so that the cost of the final optimal path is one.
@@ 8 Discussion and conclusion
Passages provide advantages over clearance in measuring accessible free space. Many approaches to enlarging path clearance, such as path search in Voronoi graphs, are unable to maximize clearance globally. Though the maximum clearance is attainable in maximum clearance planning, the degradation problem exists. Moreover, path clearance cannot be optimized as flexibly as discrete passage widths in PTOPP. Passage graphs offer a new perspective to organize objects that has potential in scene understanding. Narrow passage regions may also be used to guide sampling for efficient path finding. The Gabriel condition is proposed for passage detection because of its intuitive spatial interpretation in robotics contexts. Another possible choice of proximity graph is the relative neighborhood graph (RNG), a subgraph of the Gabriel graph (toussaint1980relative; melchert2013percolation). RNG maintains the shortest edges of Delaunay triangles, imposing a stricter passage criterion than the Gabriel condition. Beyond 2D and 3D workspace, the concept of passages is promising in higher dimensional spaces, e.g., the joint space, in describing confined configurations. But characterization of subjects, e.g., obstacles, in high dimensional space is complicated (dai2024certified).
Accessible free space optimization for paths is studied in this article as one most direct and important application of PTOPP. Other PTOPP problem instances exist since PTOPP has good configurability and applicability to formulate and solve various problems related to passages. For example, passages may be abstracted as toll stations that incur costs to pass. The principles of compatible path costs provide important guidance for cost design in OPP. The presented algorithmic framework of PTOPP is general and aligns with sampling-based optimal planners. Thus, common techniques to accelerate planning and improve the solution quality, such as bidirectional planning and biasing samples to the goal region, are applicable. Apart from sampling-based methods, pure cell-based path planning methods are enabled by Gabriel cells. They can reach a much faster planning speed with no need for a large number of random samples. Their dependent procedures, however, have been covered in the proposed algorithms.
@@ A.1 Geodesic distance traversed in Delaunay graphs in passage detection
When extracting the Gabriel graph from a Delaunay graph, it suffices to check all the edges. For edge $(\mathbf{x}_{i},\mathbf{x}_{j})$, the graph nodes $\mathbf{x}_{k}$ that possibly violate the Gabriel condition are limited to adjacent nodes of $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$. When determining passages from the Delaunay graph $\mathcal{DG}(C_{c})$ constructed from the obstacle centroids $C_{c}$, solely examining edges may miss passages because obstacles are volumetric. The solution is to increase the check range of geodesic distance. The checked geodesic distance $k_{gd}$ in $\mathcal{DG}(C_{c})$ is described in (9) and adopted as two in our implementation of (10), which proves to be adequate to find all passages under test conditions. Major factors influencing $k_{gd}$ include obstacle dimensions and distributions. The more obstacles can be treated as points (e.g., with small sizes and in large environments), the smaller $k_{gd}$ is.
Here we report detected passage data in various setups with $k_{gd}=1$ and $2$. It empirically verifies that a small $k_{gd}>1$ suffices to detect all passages. As demonstrated in Table 3 and 4, when obstacle sizes have limited variations, $k_{gd}=2$ reports all passages, e.g., for the side length range $[20,60]$ used in experiments. As expected, only checking edges of $\mathcal{DG}(C_{c})$, i.e., $k_{gd}=1$, fails to find all passages mostly. If obstacle sizes have extreme variations, $k_{gd}=2$ may also fail. When the obstacle side length range is $[1,60]$, obstacle sizes may differ by thousands of times. Failure to detect all passages occurs occasionally for $k_{gd}=2$, especially when obstacles are dense. Importantly, only a negligible number of passages are missed in failure cases as listed in Table 4. This is acceptable in most practical scenes. To find all passages, $k_{gd}$ only needs to be larger than two. $k_{gd}$ does not change the detection complexity because the expected degree of a Delaunay graph node is $O(1)$ and at most six (mark2008computational). The computational efficiency is guaranteed to outperform the brute-force check strategy.
A.2 Compatible GPW-PTOPP path cost design
In general, the cost form in (21) is not order-preserving as a sum of paired weight and width products. The design objective is to make the order of path costs consistent with the reversed lexicographic order of passage width vectors by setting appropriate weights. The lexicographic order of passage vectors is preserved in path concatenation since
where $\sqcup$ stands for inserting a new entry into one sorted vector while maintaining the sorted order. This is easily verified by comparing the new element’s insertion positions in two vectors.
Table 3: Number of trails that report all passages. For a given obstacle number, 30 trials on random obstacle distributions are conducted. The environment dimension is $1000\times 600$.
@@ A.2 Compatible GPW-PTOPP path cost design
$\mathbf{p}_{\sigma}$ is fixed sized in PTOPP and the insertion above is accomplished by first inserting the new element and then truncating the last and also the largest element. To ensure the transition in (23) for path cost (21), weight values in $\mathbf{w}_{p}$ are analyzed recursively. Start with the basic case of 2-GPW-PTOPP, i.e., $\textit{k}=2$. Let $\mathbf{p}_{\sigma_{1}}=[p_{1,1},p_{1,2}]^{\mathsf{T}},\mathbf{p}_{\sigma_{2}}=[p_{2,1},p_{2,2}]^{\mathsf{T}},\mathbf{w}_{p}=[w_{1},w_{2}]^{\mathsf{T}}$ and $\mathbf{p}_{\sigma_{1}}\prec_{\text{lex}}\mathbf{p}_{\sigma_{2}}$. $f_{c}(\sigma_{1})>f_{c}(\sigma_{2})$ implies that
$w_{1}p_{1,1}+w_{2}p_{1,2}<w_{1}p_{2,1}+w_{2}p_{2,2}.$ (31)
$p_{1,1}=p_{2,1},p_{1,2}<p_{2,2}$ is trivial. If $p_{1,1}<p_{2,1}$, the full constraint on weights is expressed as
@@
Extending to $\textit{k}=3$, a similar ratio constraint on weights is derived. If $p_{1,1}=p_{2,1}$, the constraint in (32) applies to $w_{2},w_{3}$ except that choices of passage widths decrease by one because $p_{1,1}=p_{2,1}$ is removed. But the scaled value in (33) is still applicable to set $w_{2}/w_{3}>\max\Delta p/\min\Delta p$. If $p_{1,1}<p_{2,1}$, the following inequality holds
@@
By this means, the reversed lexical orders of passage width vectors are maintained in (21). The ratios between contiguous weight items are required to be sufficiently large. In practice, any large ratio that meets (36) works. Hence, it suffices to use $w_{p,i}/w_{p,i+1}\gg 1$ in GPW-PTOPP.
There are other possible constructions of compatible GPW-PTOPP path costs, such as $f_{c}(\sigma)=-\sum p_{i}$, $p_{i}\in\mathbf{p}_{\sigma}$ and its combination with the path length $f_{c}=\text{Len}(\sigma)-\sum p_{i}$. They have simple structures and good interpretability. But the problem is that while these costs favor wide passages, they are unable to prioritize avoiding narrow passages. For example, $\mathbf{p}_{\sigma}=[1,9]^{\mathsf{T}}$ leads to the same cost as $\mathbf{p}_{\sigma}=[5,5]^{\mathsf{T}}$ though it contains a much smaller passage width. For this reason, they are not adopted in this paper for accessible free space optimization purposes.